What is the common difference (d) and the next 2 consecutive terms of the series:
1, 6/5, 7/5, 8/5, 9/5, 2, ...
Ans: d=? ; ?, ?
Ex:
15, 10, 5, 0, -5, -10
Ans: d=5; -15, -20

Answer by Cromlix(4381)

Subtract 1 from 6/5 = 1/5
Subtract 6/5 from 7/5 = 1/5
Therefore difference is 1/5
The next two consecutive terms of the series are:-
1) 2 + 1/5 = 11/5
2) 11/5 + 1/5 = 12/5
Hope this helps:-)